OpenStudy (anonymous):

Find the solutions to the systems of nonlinear equations y=2x x^2+y^2=15

OpenStudy (anonymous):

just substitute y= 2x in second eq... x^2 + (2x)^2=15 x^2 + 4x^2=5x^2=15 solve it